OpenStudy (sidney18):

A 24-foot utility pole casts a 32-foot shadow. To the nearest degree, find the angle of elevation of the sun.

OpenStudy (sidney18):

A)37° B)41° C)49° D)53°

OpenStudy (anonymous):

 \[\tan A=24/32\] evaluate for A

OpenStudy (anonymous):

it is about 37 degrees

OpenStudy (sidney18):

please explain how you got that

OpenStudy (anonymous):

 since the triangle ABC is right the following equalities are true: tanA=a/b sinA=a/c cosA=b/c

OpenStudy (sidney18):

ok thank you

OpenStudy (anonymous):

np :)